OpenStudy (anonymous):

How can I find the required sample size on a Ti83?

jimthompson5910 (jim_thompson5910):

have a look at this page https://www.ltcconline.net/greenl/courses/201/estimation/ciprop.htm let me know if it helps or not. Also, post the full problem if you can

OpenStudy (anonymous):

An IQ test is designed so that the mean is 100 and the SD is 10 for the population of normal adults. Find the sample size necessary to estimate the mean IQ score of statistics students such that it can be said with 90% confidence that the sample mean is within 2 IQ points of the true mean. Assume that the SD=10 and determine the required sample size using technology. Then determine if this is a reasonable sample size for a real world calculation.

jimthompson5910 (jim_thompson5910):

thanks

jimthompson5910 (jim_thompson5910):

did that link help at all?

jimthompson5910 (jim_thompson5910):

you'll use a formula given on that page

OpenStudy (anonymous):

I'll try it out

OpenStudy (anonymous):

Im not sure which solution to use?

OpenStudy (anonymous):

How can I put it into my ti83?

jimthompson5910 (jim_thompson5910):

you'll use the formula \[\Large n = \left(\frac{z_c*\sigma}{E}\right)^2\] as shown on that page

jimthompson5910 (jim_thompson5910):

\(\Large z_c\) is the critical z value based on the confidence level You'll find using a table or calculator that \(\Large z_c \approx 1.645\) at the 90% confidence level

jimthompson5910 (jim_thompson5910):

You are given \[\Large \sigma = 10\] \[\Large E = 2\]

OpenStudy (anonymous):

I got 16.45

jimthompson5910 (jim_thompson5910):

\[\Large n = \left(\frac{z_c*\sigma}{E}\right)^2\] \[\Large n = \left(\frac{1.645*10}{2}\right)^2\] \[\Large n = ???\]

OpenStudy (anonymous):

Yep, that is what im getting... Idk what I am doing wrong?

jimthompson5910 (jim_thompson5910):

you only computed 1.645*10 to get 16.45

jimthompson5910 (jim_thompson5910):

you forgot to divide that result by 2 then square that result

OpenStudy (anonymous):

9.069

jimthompson5910 (jim_thompson5910):

incorrect

jimthompson5910 (jim_thompson5910):

\[\Large n = \left(\frac{z_c*\sigma}{E}\right)^2\] \[\Large n = \left(\frac{1.645*10}{2}\right)^2\] \[\Large n = \left(\frac{16.45}{2}\right)^2\] \[\Large n = \left(8.225\right)^2\] \[\Large n = ???\]

OpenStudy (anonymous):

67.65

jimthompson5910 (jim_thompson5910):

then you round that up to the nearest whole number

OpenStudy (anonymous):

68

jimthompson5910 (jim_thompson5910):

so n = 68 is the required sample size

OpenStudy (anonymous):

Thank you

jimthompson5910 (jim_thompson5910):

yw